if the area of a square is 21 more the the perimeter, what is the length of a side
Answer by ankor@dixie-net.com(12704)

if the area of a square is 21 more the the perimeter, what is the length of a side
:
s^2 - 21 = 4s
:
s^2 - 4s - 21 = 0
:
(s + 3)(s - 7) = 0
:
s = +7 is the length of a side
:
Check:
7^2 - 21 = 4*7
49 - 21 = 28